OpenStudy (ayonnaleflore):

Solve the following equation by transforming it into a perfect square trinomial. x2 – 14x = 15 {–15, 29} {57, 71} {–22, 36} {–1, 15}

OpenStudy (er.mohd.amir):

for perfect sq a*a+b*b+-2a*b sue this

OpenStudy (er.mohd.amir):

x*x-14x-15=0 x*x-2*x*7+7*7-7*7-15=0

OpenStudy (calculusxy):

Make \(x^2 - 14x = 15\) into a quadratic equation in standard form (\(ax^2 + bx + c = 0\)).

OpenStudy (calculusxy):

You can bring over the 15 from the right side of the equal sign to the left side by subtracting it form both sides. \(x^2 - 14x = 15\) \(x^2 - 14x - 15 = 15 - 15\) Then you will see that on the right side, it will eventually equal to 0. \(\color{green}{x^2 - 14x - 15 = 0}\)

OpenStudy (calculusxy):

Finally use any methods, such as factoring, completing the square, quadratic formula, etc.

OpenStudy (calculusxy):

@ayonnaleflore Are you still online and solving this problem? If you don't reply, then I can't help you.